OpenStudy (anonymous):

would this be the system independent, dependent, or inconsistent? 12x +3y= 12 y+-4x=5

OpenStudy (anonymous):

3y+12x=12 divide everything by 3 and we get y+4x=12 y-4x=5 is the other equation. they are independent, they are different lines.

OpenStudy (anonymous):

they were supposed to be y+4x=4 and y-4x=5 i mean... same answer though